The reading speed of sixth grade students is approximately normal, with a mean speed of 125 words per minute and a standard deviation of 24 words per minute. a school psychologist wants to determine reading rates for unusual students (both slow and fast). Determine the reading rates that mark the minimum and maximum of the middle 95% of all sixth grade students
Answer #1

Solution:

Given mean is 125 minute
standard deviation sigma is 24 minute
Two sigmas is 48 words per minute.


Slow readers will be below 125-2(24) = 125 - 48 = 77 wpm.
Fast readers will be above 125+2(24) = 125+48 = 173 wpm.

The range for the 95% of readers is 77-173 wpm.